OpenStudy (anonymous):

If d = k and k = 7, which statement is true?

OpenStudy (anonymous):

A. d > 7 B. d < 7 C. d = 7 D. d + k = 7

OpenStudy (aravindg):

What do you think? The question is direct.

OpenStudy (anonymous):

d.

OpenStudy (aravindg):

How? o.O

OpenStudy (anonymous):

d+k=7 d=k k=7

OpenStudy (aravindg):

d=k k=7 Isn't it obvious that d=k=7??

OpenStudy (aravindg):

By your reasoning d+k=7+7=14=7? No.

OpenStudy (anonymous):

it say your wrong AravindG

OpenStudy (aravindg):

I am right. I never said d is right.

OpenStudy (haseeb96):

d=7 is correct just put K=7 into the given K equation

OpenStudy (anonymous):

d=k=7 so answer is C

geerky42 (geerky42):

yeah. transitive property of equality if a=b and b=c, then a=c It's pretty straightforward, really.

OpenStudy (anonymous):

I DID NOT PUT IN D.!!!